# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

### 12 Mar 2024 v1

@@ Abstract
${}^{1}$ept. Mathematical Sciences, Chalmers and University of Gothenburg ${}^{2}$niv. Grenoble Alpes, Inria, CNRS, Grenoble INP, Inserm U1216, CHU Grenoble Alpes, Grenoble Institut des Neurosciences, LJK ${}^{3}$niv. Lyon, CNRS, Inserm, INSA Lyon, UCBL, CREATIS Correspondence to: picchini@chalmers.se
@@ 1 Introduction
We consider simulation-based inference (SBI) for model parameters $\bm{\theta}$ in stochastic modelling. SBI refers to approaches that bypass the need to have a likelihood function $p(\bm{y}|\bm{\theta})$ (for data $\bm{y}$) available in closed form, or one that is cheap to approximate, by instead considering a generative model $f(\bm{\theta})$ or computer simulator. The simulator depends on $\bm{\theta}$ and possibly on other inputs such as covariates and a stream of pseudorandom numbers. Generating a dataset $\bm{y}^{*}$ by running the simulator $\bm{y}^{*}=f(\bm{\theta}^{*})$ at $\bm{\theta}=\bm{\theta}^{*}$ is then equivalent to simulating $\bm{y}^{*}$ from the unavailable likelihood function. In the following this simulation is written as $\bm{y}^{*}\sim p(\bm{y}|\bm{\theta}^{*})$. Established SBI methods include approximate Bayesian computation (ABC, reviewed in Sisson et al., 2018), synthetic likelihoods (Wood, 2010; Price et al., 2018), particle Markov chain Monte Carlo (Andrieu and Roberts, 2009; Andrieu et al., 2010), and several more that we discuss in Section 2. More recently, methods exploiting neural networks (NN) have gained considerable attention. Neural-based estimators have been used to produce approximated likelihoods or posterior distributions, typically using (discrete) normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021). NN have been used to approximate the likelihood function in implicit models (Papamakarios et al., 2019; Chen et al., 2021), the posterior distribution (Papamakarios and Murray, 2016; Greenberg et al., 2019; Durkan et al., 2020; Chen et al., 2021) or the likelihood and the posterior distribution simultaneously (Wiqvist et al., 2021; Radev et al., 2023a).
@@
These neural-based posterior and likelihood estimation methods have produced exceptionally flexible inference strategies for models with intractable likelihoods. In this work, we propose to investigate more frugal strategies (Evchenko et al., 2021) that can run with limited resources and a much smaller computational footprint, while returning inference of similar quality as neural-based approaches. For SBI, we propose to investigate structured mixtures of probability distributions whose versatility has been widely recognized for a variety of data and tasks, while not requiring excessive design effort or tuning. Their expressivity makes them good candidates to account for complex multivariate models both for the likelihood and posterior distribution. Moreover, mixture models have a much smaller number of hyperparameters compared to NN, and this makes them more amenable to interpretation and efficient learning (see Table 1). More specifically, we design a new simulation-based inference method named Sequential Mixture Posterior and Likelihood Estimation (SeMPLE). SeMPLE learns approximations for both the posterior $p(\bm{\theta}|\bm{y})$ and the likelihood $p(\bm{y}|\bm{\theta})$ using a structured mixture model obtained via the Gaussian Locally Linear Mapping (GLLiM, Deleforge et al., 2014) method. Interestingly, both approximations are jointly returned following a run of an Expectation-Maximization (EM) algorithm within GLLiM. We then use the GLLiM approximate posterior as a well-informed proposal distribution in a tuning-free Metropolis-Hastings sampler. The procedure is embedded in a sequential strategy, which we compare with state-of-the-art NN-based sequential algorithms for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Other algorithms could be considered for comparison, however we stick with SNL and SNPE-C and a motivation for this is given in Section 2. Our comparisons are in terms of inference quality as well as resource and energy requirements. In general, we did not compare against ABC samplers such as SMC-ABC, except for the Lotka-Volterra example where such comparison was more suitable, as in Lueckmann et al., 2021 it was shown that standard implementations of ABC approaches usually require a larger number of model simulations, compared to NN-based approaches, although recent research improved the SMC-ABC performance (Picchini and Tamborrino, 2022). We found that, for a given computational budget specified by the total number of model simulations, our method is cheaper to run and produces accurate inference thanks to (i) the informed proposal sampler created by GLLiM and (ii) the computationally faster model training. As an illustration, Figure 1 shows inference for three datasets of the Two Moons model, a model with bimodal posterior that we further analyse in Section 5. The runtime, energy and memory requirements for SeMPLE are considerably lower, see Table 1 and Figure 2. Overall SeMPLE shows low resources costs that are always of the same magnitude ($<10$ kJ and $<1$ GB), while the requirements for SNL and SNPE-C vary much more with the model dimension or the inference complexity. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance.
@@
Figure 1: Two Moons model from SBIBM, datasets 1 (top), 8 (middle) and 9 (bottom). Posterior samples after 4 rounds of SeMPLE (column 1), 10 rounds of SNPE-C and SNL (columns 2 and 3), using 10K model simulations. The true posteriors are in column 4.
@@
Table 1: Median resource requirements (across 10 independent repetitions) and number of model hyperparameters $\phi$ for posterior inference with a budget of 10K (two moons and Bernoulli GLM) and 40K (hyperboloid and Ornstein-Uhlenbeck) model simulations. Best (lowest) values are in bold. SNPE-C and SNL used the default SBIBM setup.
@@
Section 2 gives a brief overview of the main SBI approaches. Section 3 describes how GLLiM can be used to construct surrogate posteriors and surrogate likelihoods, while the complete SeMPLE methodology is explained in Section 4. Results from SeMPLE, SNL and SNPE-C are then illustrated in Section 5. Test models, partly reported in the Supplementary Material, include six simulation studies comprising multimodal surfaces (two moons, multiple hyperboloid), stochastic differential equations (Ornstein-Uhlenbeck), generalised linear models (Bernoulli GLM), a Markov jump process with additional measurement error (Lotka-Volterra), and highly correlated posteriors with 20 parameters (twisted-prior). Code is available at https://github.com/henhagg/semple.
@@ 2 Related work
Simulation-based inference, also referred to as likelihood-free inference, is reviewed e.g. in Cranmer et al. (2020) and Pesonen et al. (2023). In a Bayesian framework the focus is on posterior estimation, which in SBI directly relates to the quality of the approximations made to the unavailable likelihood, the posterior distribution or both. This aspect raises the issue of how to evaluate the quality of posterior inference when a reference ground truth posterior is absent by definition. This is still an open question and is further discussed in Section 6. A second important ingredient is sampling efficiency, as simulation-based methods are typically relying on exploration, and on Monte Carlo principles that can be prohibitively expensive. Active or sequential learning is a way to address sample efficiency by gradually discovering and guiding where simulations should be made for a specific given observation. Our proposed method is sequential and we focus on comparisons with NN-based sequential methods for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Many other approaches exist for Bayesian inference with intractable likelihoods, such as sequential neural ratio estimation (SNRE, Hermans et al., 2020), the several variants of ABC such as SMC-ABC (Toni et al., 2009; Del Moral et al., 2012), synthetic likelihoods (Wood, 2010; Price et al., 2018) and particle MCMC (Andrieu and Roberts, 2009; Andrieu et al., 2010). Comparisons between several of the above mentioned methods are available in previous works. For example, Greenberg et al. (2019) show that SNPE-C is computationally more efficient than the predecessors [^1] The acronyms SNPE-A and SNPE-B follow the taxonomy in Papamakarios et al. (2019). SNPE-A (Papamakarios and Murray, 2016) and SNPE-B (Lueckmann et al., 2017), where efficient means that, for a given computational budget, SNPE-C approaches the true posterior using a smaller number of model simulations. Lueckmann et al. (2021) show that, for nontrivial case-studies, both SNPE-C and SNRE are more efficient than SNL and SMC-ABC. Moreover, SNPE-C and SNRE perform similarly. The above motivates why in this work we focus on the comparison with SNPE-C using the SBIBM package implementation (Lueckmann et al., 2021) (an exception is the Lotka-Volterra Markov jump process, where we compare against SMC-ABC, as an objective comparison with NN-based methods is difficult here, see section 5.3 for more details). Since SeMPLE includes an MCMC step, same as SNL, it is also relevant to compare against SNL, using its implementation in SBIBM, which is an ameliorated one, compared to the original version of Papamakarios et al. (2019).
@@
Note that there are similarities between our SeMPLE and SNPE-A, as detailed in Supplementary Section 3. Briefly, in SNPE-A a $K$ component mixture density network (MDN) is used as a proposal sampler. As such, the MDN means and covariances are parametrised via neural networks. However SNPE-A does not have a built-in strategy to learn each of the $K$ components. Instead, the GLLiM procedure embedded into SeMPLE automatically determines the means and covariance matrices for each of the $K$ components, thus increasing flexibility to the shapes of the posteriors that can be targeted. The fact that SeMPLE does not use MDNs is key to its resource-efficiency, since no complex NN architectures are employed in SeMPLE, and the benefit of the EM algorithm can be exploited towards fitting the mixture models.
@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
The modelling power of GLLiM comes essentially from the fact that both conditional distributions above are mixtures of Gaussian affine experts (MoE), which are known for their modelling and approximation capabilities (Nguyen et al., 2019, 2021, 2023). As such, they are good candidates to provide surrogate likelihoods and posteriors. All the more so as, given a training dataset made of many pairs of $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$, surrogate models specified via GLLiM can be estimated via an EM algorithm, see Deleforge et al. (2014) and Forbes et al. (2022) for details. In fact a surrogate likelihood and posterior can be obtained with $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$ prior-predictive realizations simulated from $p(\bm{\theta})p(\bm{y}\,|\,\bm{\theta})$, but more generally, as exploited in SeMPLE, GLLiM can be used in a sequential way, by changing the training set to include posterior draws.
@@
3:     for $n=1:N$ do
4:         if $r=0$ or $r=1$ sample $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ without MCMC, otherwise sample $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ using Algorithm 2.
5:         Sample (implicitly) $\bm{y}_{n}\sim p(\bm{y}\,|\,\bm{\theta}_{n})$.
6:     end for
7:     if r = 0 then
8:         Collect $\mathcal{D}_{0}=\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$.
9:         Train $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{0}$. Update $K$.
10:         Obtain $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$.
11:         Set $\tilde{p}_{1}(\bm{\theta})\leftarrow q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}=\bm{y}_{o})$.
12:     else
13:         Collect $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}\setminus\mathcal{D}_{0}$.
14:         Train $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{r}$. Update $K$.
15:         Obtain $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y})$.
16:         Set $\tilde{p}_{r+1}(\bm{\theta})\leftarrow q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$  ((optional) $\tilde{p}_{r+1}(\bm{\theta})\leftarrow\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$)
17:     end if
18:end for

### 16 Mar 2024 v2

@@ 1 Introduction
We consider simulation-based inference (SBI) for model parameters $\bm{\theta}$ in stochastic modelling. SBI refers to approaches that bypass the need to have a likelihood function $p(\bm{y}|\bm{\theta})$ (for data $\bm{y}$) available in closed form, or one that is cheap to approximate, by instead considering a generative model $f(\bm{\theta})$ or computer simulator. The simulator depends on $\bm{\theta}$ and possibly on other inputs such as covariates and a stream of pseudorandom numbers. Generating a dataset $\bm{y}^{*}$ by running the simulator $\bm{y}^{*}=f(\bm{\theta}^{*})$ at $\bm{\theta}=\bm{\theta}^{*}$ is then equivalent to simulating $\bm{y}^{*}$ from the unavailable likelihood function. In the following this simulation is written as $\bm{y}^{*}\sim p(\bm{y}|\bm{\theta}^{*})$. Established SBI methods include approximate Bayesian computation (ABC, reviewed in Sisson et al., 2018), synthetic likelihoods (Wood, 2010; Price et al., 2018), particle Markov chain Monte Carlo (Andrieu & Roberts, 2009; Andrieu et al., 2010), and several more that we discuss in Section 2. More recently, methods exploiting neural networks (NN) have gained considerable attention. Neural-based estimators have been used to produce approximated likelihoods or posterior distributions, typically using (discrete) normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2021). NN have been used to approximate the likelihood function in implicit models (Papamakarios et al., 2019; Chen et al., 2021), the posterior distribution (Papamakarios & Murray, 2016; Greenberg et al., 2019; Durkan et al., 2020; Chen et al., 2021) or the likelihood and the posterior distribution simultaneously (Wiqvist et al., 2021; Radev et al., 2023a).
@@
These neural-based posterior and likelihood estimation methods have produced exceptionally flexible inference strategies for models with intractable likelihoods. In this work, we propose to investigate more frugal strategies (Evchenko et al., 2021) that can run with limited resources and a much smaller computational footprint, while returning inference of similar quality as neural-based approaches. For SBI, we propose to investigate structured mixtures of probability distributions whose versatility has been widely recognized for a variety of data and tasks, while not requiring excessive design effort or tuning. Their expressivity makes them good candidates to account for complex multivariate models both for the likelihood and posterior distribution. Moreover, mixture models have a much smaller number of hyperparameters compared to NN, and this makes them more amenable to interpretation and efficient learning (see Table 1). More specifically, we design a new simulation-based inference method named Sequential Mixture Posterior and Likelihood Estimation (SeMPLE). SeMPLE learns approximations for both the posterior $p(\bm{\theta}|\bm{y})$ and the likelihood $p(\bm{y}|\bm{\theta})$ using a structured mixture model obtained via the Gaussian Locally Linear Mapping (GLLiM, Deleforge et al., 2014) method. Interestingly, both approximations are jointly returned following a run of an Expectation-Maximization (EM) algorithm within GLLiM. We then use the GLLiM approximate posterior as a well-informed proposal distribution in a tuning-free Metropolis-Hastings sampler. The procedure is embedded in a sequential strategy, which we compare with state-of-the-art NN-based sequential algorithms for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Other algorithms could be considered for comparison, however we stick with SNL and SNPE-C and a motivation for this is given in Section 2. Our comparisons are in terms of inference quality as well as resource and energy requirements. In general, we did not compare against ABC samplers such as SMC-ABC, except for the Lotka-Volterra example where such comparison was more suitable, as in Lueckmann et al., 2021 it was shown that standard implementations of ABC approaches usually require a larger number of model simulations, compared to NN-based approaches, although recent research improved the SMC-ABC performance (Picchini & Tamborrino, 2022). We found that, for a given computational budget specified by the total number of model simulations, our method is cheaper to run and produces accurate inference thanks to (i) the informed proposal sampler created by GLLiM and (ii) the computationally faster model training. As an illustration, Figure 1 shows inference for three datasets of the Two Moons model, a model with bimodal posterior that we further analyse in Section 5. The runtime, energy and memory requirements for SeMPLE are considerably lower, see Table 1 and Figure 2. In other examples, (e.g. Figure 3), the speedup provided by SeMPLE is even more dramatic. Overall SeMPLE shows low resources costs that are always of the same magnitude ($<10$ kJ and $<1$ GB), while the requirements for SNL and SNPE-C vary much more with the model dimension or the inference complexity. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance.
@@
Figure 1: Two Moons model from SBIBM, datasets #1 (top), #8 (middle) and #9 (bottom). Posterior samples after 4 rounds of SeMPLE (column 1), 10 rounds of SNPE-C and SNL (columns 2 and 3), using 10K model simulations. The true posteriors are in column 4.
Figure 2: Two Moons. (left) Median C2ST (the lower the better) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
@@
Table 1: Median resource requirements (across 10 independent repetitions) and number of model hyperparameters $\phi$ for posterior inference with a budget of 10K (two moons and Bernoulli GLM) and 40K (hyperboloid and Ornstein-Uhlenbeck) model simulations. Best (lowest) values are in bold. SNPE-C and SNL used the default SBIBM setup described in Appendix A.
@@
Section 2 gives a brief overview of the main SBI approaches. Section 3 describes how GLLiM can be used to construct surrogate posteriors and surrogate likelihoods, while the complete SeMPLE methodology is explained in Section 4. Results from SeMPLE, SNL and SNPE-C are then illustrated in Section 5. Test models, partly reported in the Appendices, include six simulation studies comprising multimodal surfaces (two moons, multiple hyperboloid), stochastic differential equations (Ornstein-Uhlenbeck), generalised linear models (Bernoulli GLM), a Markov jump process with additional measurement error (Lotka-Volterra), and highly correlated posteriors with 20 parameters (twisted-prior). Code is available at https://github.com/henhagg/semple.
@@ 2 Related work
Simulation-based inference, also referred to as likelihood-free inference, is reviewed e.g. in Cranmer et al. (2020) and Pesonen et al. (2023). In a Bayesian framework the focus is on posterior estimation, which in SBI directly relates to the quality of the approximations made to the unavailable likelihood, the posterior distribution or both. This aspect raises the issue of how to evaluate the quality of posterior inference when a reference ground truth posterior is absent by definition. This is still an open question and is further discussed in Section 6. A second important ingredient is sampling efficiency, as simulation-based methods are typically relying on exploration, and on Monte Carlo principles that can be prohibitively expensive. Active or sequential learning is a way to address sample efficiency by gradually discovering and guiding where simulations should be made for a specific given observation. Our proposed method is sequential and we focus on comparisons with NN-based sequential methods for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Many other approaches exist for Bayesian inference with intractable likelihoods, such as sequential neural ratio estimation (SNRE, Hermans et al., 2020), the several variants of ABC such as SMC-ABC (Toni et al., 2009; Del Moral et al., 2012), synthetic likelihoods (Wood, 2010; Price et al., 2018) and particle MCMC (Andrieu & Roberts, 2009; Andrieu et al., 2010). Comparisons between several of the above mentioned methods are available in previous works. For example, Greenberg et al. (2019) show that SNPE-C is computationally more efficient than the predecessors [^1] The acronyms SNPE-A and SNPE-B follow the taxonomy in Papamakarios et al. (2019). SNPE-A (Papamakarios & Murray, 2016) and SNPE-B (Lueckmann et al., 2017), where efficient means that, for a given computational budget, SNPE-C approaches the true posterior using a smaller number of model simulations. Lueckmann et al. (2021) show that, for nontrivial case-studies, both SNPE-C and SNRE are more efficient than SNL and SMC-ABC. Moreover, SNPE-C and SNRE perform similarly. The above motivates why in this work we focus on the comparison with SNPE-C using the SBIBM package implementation (Lueckmann et al., 2021) (an exception is the Lotka-Volterra Markov jump process, where we compare against SMC-ABC, as an objective comparison with NN-based methods is difficult here, see section 5.3 for more details). Since SeMPLE includes an MCMC step, same as SNL, it is also relevant to compare against SNL, using its implementation in SBIBM, which is an ameliorated one, compared to the original version of Papamakarios et al. (2019).
@@
Note that there are similarities between our SeMPLE and SNPE-A, as detailed in Appendix C. Briefly, in SNPE-A a $K$ component mixture density network (MDN) is used as a proposal sampler. As such, the MDN means and covariances are parametrised via neural networks. However SNPE-A does not have a built-in strategy to learn each of the $K$ components. Instead, the GLLiM procedure embedded into SeMPLE automatically determines the means and covariance matrices for each of the $K$ components, thus increasing flexibility to the shapes of the posteriors that can be targeted. The fact that SeMPLE does not use MDNs is key to its resource-efficiency, since no complex NN architectures are employed in SeMPLE, and the benefit of the EM algorithm can be exploited towards fitting the mixture models.
@@ 3 Amortized surrogates for likelihoods and posteriors via GLLiM
The modelling power of GLLiM comes essentially from the fact that both conditional distributions above are mixtures of Gaussian affine experts (MoE), which are known for their modelling and approximation capabilities (Nguyen et al., 2019; 2021; 2023). As such, they are good candidates to provide surrogate likelihoods and posteriors. All the more so as, given a training dataset made of many pairs of $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$, surrogate models specified via GLLiM can be estimated via an EM algorithm, see Deleforge et al. (2014) and Forbes et al. (2022) for details. In fact a surrogate likelihood and posterior can be obtained with $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$ prior-predictive realizations simulated from $p(\bm{\theta})p(\bm{y}\,|\,\bm{\theta})$, but more generally, as exploited in SeMPLE, GLLiM can be used in a sequential way, by changing the training set to include posterior draws.
@@
3:     if $r=0$ or $r=1$ sample $\{\theta_{n}\}_{n=1}^{N}$ iid $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ (without MCMC), $n=1,...,N$.
4:     Otherwise for $r\geq 2$ sample $\{\bm{\theta}_{n}\}_{n=1}^{N}$ via $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ using Algorithm 2.
5:     Conditionally on every $\bm{\theta}_{n}$, obtain $\{\bm{y}_{n}\}_{n=1}^{N}$ by sampling (implicitly) $\bm{y}_{n}\sim p(\bm{y}\,|\,\bm{\theta}_{n})$ ($n=1,...,N$).
6:     if r = 0 then
7:         Collect $\mathcal{D}_{0}=\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$.
8:         Train $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{0}$. Update $K$.
9:         Obtain $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$.
10:         Set $\tilde{p}_{1}(\bm{\theta})\leftarrow q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}=\bm{y}_{o})$.
11:     else
12:         Collect $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}\setminus\mathcal{D}_{0}$.
13:         Train $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{r}$. Update $K$.
14:         Obtain $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y})$.
15:         Set $\tilde{p}_{r+1}(\bm{\theta})\leftarrow q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$  ((optional) $\tilde{p}_{r+1}(\bm{\theta})\leftarrow\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$)
16:     end if
17:end for